OpenStudy (anonymous):

Find length of curve x = ln (t), y = sqrt(t+1) 1<=t<=2

OpenStudy (anonymous):

first you differentiate y and x

OpenStudy (anonymous):

and then put it in this for square root of (1+(dx?dy)^2)

OpenStudy (anonymous):

and do the integral since you are given the boundries

OpenStudy (anonymous):

good try but not quite

OpenStudy (anonymous):

We're actually dealing with parametric curves at this stage, so I assume it's the sqrt of (dx/dt)^2 + (dy/dt)^2 dt. I can't seem to simplify the algebra.

OpenStudy (anonymous):

do you like algebra

OpenStudy (anonymous):

Yes, for the most part. I suspect that I have done something wrong because the answer looks worse than the original problem.

OpenStudy (anonymous):

you r right

OpenStudy (anonymous):

i have 2 go kay kay

OpenStudy (anonymous):

well it was the same whether use the equation but Kix001 equation is most accurate because it a parametraization

OpenStudy (turingtest):

(dx/dt)^2=1/t^2 (dy/dt)^2=1/(t+1) \[\int\limits_{1}^{2}{\sqrt{t^2+t+1\over t^2(t+1)}}dt\]seems to be where I'm stuck at...

OpenStudy (anonymous):

\[\int_{1}^{2} \sqrt{((\log{t})\prime)^2+((\sqrt{t+1)}\prime)^{2}}\] so \[\int_{0}^{1} \sqrt{\frac{1}{t^2}+\frac{1}{(2\sqrt{t+1)^2}}}\] \[\int_{0}^{1} \sqrt{\frac{1}{t^2}+\frac{1}{4t+1}}\] \[\int_{0}^{1} \sqrt{\frac{t^2+4t+1}{4t^4+t^2}} =0.680816 \;\;\; Aprox.\]

OpenStudy (anonymous):

Made a mistake, never mind the above.

OpenStudy (turingtest):

I did lose a two in my answer, so I'm at\[\int\limits\limits_{1}^{2}{\sqrt{t^2+4t+4\over4t^2(t+1)}}dt=\int\limits\limits_{1}^{2}{(t+2) \over2t \sqrt{t+1}}dt\]umm...

OpenStudy (anonymous):

Same here. I have the same answer, but don't know how to solve it further.

OpenStudy (anonymous):

same as me

OpenStudy (turingtest):

Maybe\[\int\limits_{1}^{2}{1\over2\sqrt{t+1}}+{1\over t \sqrt{t+1}}dt\]now about that second term... I'm thinking either integration by parts, or it might be a form in some table. Let me check...

OpenStudy (anonymous):

but that is how did you get the second term

OpenStudy (turingtest):

how? by what?

OpenStudy (turingtest):

http://www.wolframalpha.com/input/?i=integral+1%2F%28t+sqrt%28t%2B1%29%29 but I can't get it to show steps...

OpenStudy (turingtest):

oh, you asked how I split up the fraction?

OpenStudy (turingtest):

anyway, Wolframs answer works, but it involves two substitutions.\[\sqrt{t+1}+\ln(1-\sqrt{t+1})-\ln(1+\sqrt{t+1})\]evaluated from 1 to 2 is the answer.

OpenStudy (anonymous):

Ouch, did you get that via integration by parts?

OpenStudy (turingtest):

go to Wolfram on the link and look at "show steps" in the upper right: http://www.wolframalpha.com/input/?i=integral+1%2F%28t+sqrt%28t%2B1%29%29 Like I said, two tricky substitutions were used here. There's probably another way.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

if you are here let me know